# Probing soft signals of gravitational-wave memory with space-based interferometers

### 30 Mar 2026 v1

@@ III.2 Displacement memory from compact binary scattering
$\frac{f\tilde{h}_{2,0}}{h_{0}}\approx$ $\sqrt{\frac{2}{3}}\,\varpi\ln 2e\varpi.$
The correction factor depends on $\varpi$ and $e$. It turns out that the LO correction to $\text{Im}\,\tilde{h}_{\lambda}$ is linear, with $a_{0}=\pi^{2}/\Omega$; while the LO correction to $\text{Re}\,\tilde{h}_{\lambda}$ is logarithmic. These corrections are associated with the $1/t$ decay of the waveform as $t\to\pm\infty$ (see Sec. II). As a measure of the relative deviation from the soft waveform, we show the contours of $|f\tilde{h}_{2,\pm 2}/(\lim_{f\to 0}f\tilde{h}_{2,\pm 2})-1|$ and $|f\tilde{h}_{2,0}/(\lim_{f\to 0}f\tilde{h}_{2,\pm 2})|$ in Fig. 1. The deviations decrease with $e$ in the shown parameter space, and become less than 0.1 at $f\lesssim 10^{-2}\Omega/\pi$. The deviation of $\text{arg}\,\tilde{h}$ at given $\{\Theta,\Phi\}$ from the soft limit is smaller (see Fig. 2 for a concrete example). Therefore, a soft waveform with a real correction factor $C(f)$ provides a good approximation to $\tilde{h}(f)$ at sufficiently low frequencies.
From Eq. (25), the memory amplitude $|\Delta h|=\sqrt{\Delta h_{+}^{2}+\Delta h_{\times}^{2}}$ is given by
$\frac{|\Delta h|}{\frac{\nu M}{R}\frac{\sqrt{e^{2}-1}}{ae^{2}}}=\sqrt{-8s^{4}_{\Theta}c_{4\Phi}+28c_{2\Theta}+c_{4\Theta}+35},$ (26)
thus $|\Delta h|$ is proportional to the quantity:
$\frac{M\sqrt{e^{2}-1}}{ae^{2}}=\frac{b/M}{(b/M)^{2}+1/v_{0}^{4}},$ (27)
which increases with $v_{0}$ for given $b$, and peaks at $b=M/v_{0}^{2}$ for given $v_{0}$. For given $b$ and $v_{0}$, $|\Delta h|$ takes the maximum value $8$ at $\Theta\in\{0,\pi\}$, and vanishes at $\Theta=\pi/2\,\land\,\Phi\,\text{mod}\,\pi/2=0$.
@@
$\lim_{n\to\infty}H_{in}$ $\to-\frac{ie^{-n\left(\sqrt{e^{2}-1}-\text{arcsec}\,e\right)}}{(e^{2}-1)^{1/4}\sqrt{n\pi/2}},$ (28)
$\lim_{n\to\infty}H_{1+in}$ $\to-\frac{\left(\sqrt{e^{2}-1}+i\right)e^{-n\left(\sqrt{e^{2}-1}-\text{arcsec}\,e\right)}}{e(e^{2}-1)^{1/4}\sqrt{n\pi/2}}.$
@@
Figure 2: Relative deviations from the soft waveform in $\text{arg}\,\tilde{h}$ for various eccentricities, at $\{\Theta,\Phi\}=\{\pi/6,\pi/3\}$.
Many of the features outlined above persist once PN corrections are included, at least when those corrections are not too large. We include the PN corrections to the CM-frame equation of motion (EOM) $\ddot{\mathbf{x}}=-(M/r^{3})\,\mathbf{x}+\mathbf{F}$ and the gravitational waveform according to [60]. The initial condition is set by an osculating hyperbolic orbit with orbital elements $\{a,e\}$ and the total mass parameter $M$ (see, e.g., [15]). Such a choice for the osculating orbit is not unique, since the initial orbit depends only on $\mathbf{v}_{0}$ and $\mathbf{b}$. However, this setting enable us to directly compare the results of Newtonian and PN EOMs in the harmonic coordinates. We choose the initial conditions such that the osculating Keplerian orbit coincides with the one considered above. For simplicity, we neglect the spins of the two bodies, thus $\mathbf{F}=\mathbf{F}_{\text{1PN}}+\mathbf{F}_{\text{2PN}}+\mathbf{F}_{\text{2.5PN}}+\mathcal{O}(\text{3PN})$; the PN orbit also lies in the $XY$ plane, but its periastron deviates from that of the Keplerian orbit. At 1PN order, the change of the argument of periastron $\varphi_{0}$ of the osculating orbit during the scattering is given by [15, 11]
@@
For concreteness, we choose $\nu=0.2$, $b/M=100$ and $v_{0}=0.13$, the results are shown in Fig. 3. To verify the convergence of PN corrections, we perform the computation at different PN orders. The approximation (29) (shown as the thick gray line in the top panel of Fig. 3) is found to agree well with the 2.5PN orbit, indicating that higher-order PN corrections are relatively small. The orbit is therefore well described by the 1PN-accurate quasi-Keplerian parameterization [23]. Nevertheless, there is a considerable difference between the 0PN and 0.5PN waveforms, but the 1PN correction to the waveform is already small. It should be noted that the null memory contribution to the waveform in the present case enters at 2.5PN order [32], and is thus negligible.
@@
As a test, we verify that the spectrum of 0PN waveform computed using 0PN orbit agrees perfectly with Eqs. (22)-(23). The spectra of $|\text{Re}\,\tilde{h}_{\pm}(f)|$ and $|\text{Im}\,\tilde{h}_{\pm}(f)|$ of the 1PN waveforms are shown in the bottom panel of Fig. 3. The numeric spectrum at the lowest frequency is pretty close to the soft waveform with $\Delta h_{\lambda}$ computed in the time domain. The spectrum of $f\,\text{Im}\,\tilde{h}_{\pm}$ at low frequencies is well described by a linear correction: $f\,\text{Im}\,\tilde{h}_{\lambda}\propto 1+\kappa_{\lambda}\varpi$, with $\kappa_{+}\approx 0.95\pi$ and $\kappa_{\times}\approx 0.92\pi$. These slightly deviate from the Keplerian value $\kappa=\pi$ when expanded near $f=0$, as given by Eq. (25). Notably, the correction factors of $f\,\text{Im}\,\tilde{h}_{+}(f)$ and $f\,\text{Im}\,\tilde{h}_{\times}(f)$ are very close to each other at all frequencies. As the frequency increases, the real part of the spectrum becomes more important.
At sufficiently high frequency, $\text{Im}\,\tilde{h}_{+}\approx\text{Re}\,\tilde{h}_{\times}$ and $\text{Re}\,\tilde{h}_{+}\approx\text{Im}\,\tilde{h}_{\times}$, indicating that the radiation is almost purely right-handed polarized. This is also reflected in the 0PN spectra of the $Z$-component angular momentum and graviton number fluxes, whose ratio approaches 2 at high frequencies, as for an elliptical orbit. The decaying behavior at high frequencies is similar to the 0PN approximation given by Eq. (28), as depicted by the green lines in the bottom panel of Fig. 3.
@@
Fig. 4 shows the spectrum $\tilde{h}^{\prime}(f)$ obtained by an overall phase rotation that transforms $\tilde{h}$ at the lowest frequency of the numeric spectrum to a purely imaginary value. We see that $\tilde{h}^{\prime}(f)$ is well described by a soft waveform with a real correction factor $C(f)=1+\kappa\varpi+\mathcal{O}(\varpi^{2})$ at low frequencies. In this example, $\kappa\approx 2.63\pi$, which is quite different from $\kappa_{\lambda}$ describing the linear correction to $f\,\text{Im}\,\tilde{h}_{\lambda}$.
@@
Figure 3: CM-frame orbits (top panel) and gravitational waveforms in the time domain (middle panel) and frequency domain (bottom panel) of a compact binary scattering, with $\nu=0.2$, $b/M=100$, $v_{0}=0.13$ (corresponding to $e=2$), and $\{\Theta,\Phi\}=\{\pi/6,\pi/3\}$. In the middle panel, the 0PN waveform is computed using the 0PN orbit, whereas the 0.5PN and 1PN waveforms are computed using the 2PN orbit. The bottom panel shows the spectrum of the 1PN waveform, with $\Omega=\sqrt{M/a^{3}}=v_{0}^{3}/M\approx\pi(v_{0}/0.13)^{3}(100\,M_{\odot}/M)\times 1.4\,\text{Hz}$. Before computing the spectrum, we shift the time coordinate such that the periastron crossing happens at $t=0$. The horizontal line in the bottom panel represents the soft waveform corresponding to $\Delta h_{\lambda}$ in the time domain.
@@
Figure 4: Spectra of $f\,\text{Im}\,\tilde{h}^{\prime}(f)$ and $f\,\text{Re}\,\tilde{h}^{\prime}(f)$. The linear and quadratic fits to the low-frequency part of the spectra are also shown. The horizontal line represents the soft waveform corresponding to $\Delta h_{\lambda}$ in the time domain.
@@
For an extreme mass-ratio scattering involving a large black hole (BH), the orbit of the small compact body can be approximated by an unbound geodesic in the background BH spacetime. The dissipative effect is expected to be negligible in this case, since the incident velocity for scattering is bounded from below by the critical velocity for plunge in the test-body approximation. We compute the ordinary displacement memory produced by such a system using the general formula [10]:
@@
This is derived from the linearized Einstein equation in a flat background, but can be applied to scattering processes provided that the initial and final states live in a flat background. Eq. (30) can also be derived from the soft graviton theorem [78]. As can be checked, the low-velocity expansion of Eq. (30) up to $\mathcal{O}(v^{4})$ agrees with the 1PN approximation:
@@
$\quad+(1-3\nu)\left[(\mathbf{v}\cdot\hat{\mathbf{k}})^{2}+v^{2}/2\right]\Big)\Big]\frac{2\nu M}{R},$
for $m_{2}\ll m_{1}$. For simplicity, we consider a non-spinning central BH, and thus compute the gauge-invariant deflection angle $\chi$ of the small test body with mass $\nu M$ in a Schwarzschild spacetime with mass parameter $M$. The source frame is again chosen such that the initial orbit lies in the $XY$ plane, the initial and final three-velocities are along the direction $\mathbf{N}_{\text{in}}=c_{\vartheta}\mathbf{e}_{X}+s_{\vartheta}\mathbf{e}_{Y}$ and $\mathbf{N}_{\text{out}}=-c_{\vartheta}\mathbf{e}_{X}+s_{\vartheta}\mathbf{e}_{Y}$, respectively, with $\vartheta=(\chi-\pi)/2$. The displacement memory from Eq. (30) is given by
@@
Figure 5: Ordinary displacement memory from the scattering of a test body with mass $\nu M$ by a non-spinning BH with mass $M$, for $b/M=100$, and $\{\Theta,\Phi\}=\{\pi/5,\pi/7\}$. The blue and black solid lines show the test-body approximation. The red (green) dashed lines show the 0PN (1PN) approximations. The pink dashed lines show the results computed from Eq. (32), using the 1PN approximation to the deflection angle. The critical incident velocity $v_{\text{plunge}}\approx 0.04$. The figure inset shows the results for small incident velocities, with the horizontal axis given by $v_{0}-v_{\text{plunge}}$. For comparison, the gray dot-dashed lines show the results for a central BH with dimensionless spin $\chi=1$ along the $Z$-axis, for which $v_{\text{plunge}}=0.02$.
@@
The deflection angle depends on the initial velocity $v_{0}$ and the dimensionless impact parameter $b/M$. For the small body to be scattered rather than plunge into the BH, the impact parameter needs to exceed a critical value: ${b_{\text{plunge}}}/{M}=\sqrt{4+\frac{\sqrt{8v_{0}^{2}+1}-1}{2v_{0}^{4}}+\frac{4\sqrt{8v_{0}^{2}+1}+10}{v_{0}^{2}}}$. For a null particle, $b_{\text{plunge}}(v_{0}=1)=3\sqrt{3}M$. Likewise for a given impact parameter, $v_{0}$ needs to exceed a critical value $v_{\text{plunge}}$ in order to escape from the BH. It is possible for high-energy massless or ultra-relativistic particles (such as photons and neutrinos) to be scattered by a BH, but this cannot produce an observable memory, since even a photon with $\omega=10\,\text{TeV}$ amounts to a mass of only $9\times 10^{-54}\,M_{\odot}$.
@@
For the scattering of timelike particles, a concrete example is shown in Fig. 5, where the results are also compared with the 0PN approximation in Eq. (25) (red dashed lines) and the results computed using Eq. (31) and the 1PN approximation to the deflection angle: $\chi=2\arccos(-1/e)-\pi+\Delta\varphi_{0}$ (green dashed lines), with $\Delta\varphi_{0}$ given by Eq. (29). For comparison, we also show the results (pink dashed lines) computed from Eq. (32), but using the 1PN approximation to the deflection angle, which remains in good agreement with the test-body approximation even at high velocities. As can be seen, the 0PN approximation is accurate only for intermediate values of $v_{0}$, and the situation near $v_{\text{plunge}}$ cannot be captured by the 1PN deflection angle either, where the produced memory is relatively small. For $v_{0}\gtrsim v_{\text{plunge}}$, the periastron occurs at $r\gtrsim 4M$; the result is therefore sensitive to the spin of the central BH. Eq. (32) can also be applied to the scattering orbit in the ecliptic plane of Kerr spacetime; the gray dot-dashed lines in Fig. 5 show the result for a central BH with dimensionless spin $\chi=1$ along the $Z$-axis, which are, interestingly, closer to the pink dashed lines. At high velocities, the amplitudes of memory are considerably larger than the 0PN prediction. As in the 0PN approximation, $|\Delta h|$ is invariant under the transformation $\Theta\to\pi-\Theta$ or $\Phi\to\pi-\Phi$, but its extreme point deviates from $\Theta\in\{0,\pi\}$. For $\Theta=\pi/2$, $|\Delta h|$ still vanishes at four values of $\Phi$, two of which occur at $\pi/2$ and $3\pi/2$, while the remaining two are displaced from 0 and $\pi$.
III.3 Displacement memory from quasi-circular non-precessing BBH merger

### 13 Apr 2026 v2

@@ III.2 Displacement memory from compact binary scattering
$\frac{f\tilde{h}_{2,0}(f)}{h_{0}}\approx$ $\sqrt{\frac{2}{3}}\,\varpi\ln 2e\varpi.$
It turns out that the LO correction to $\text{Im}\,\tilde{h}_{\lambda}$ is linear, with correction factor $1+(\pi^{2}/\Omega)f$, whereas the LO correction to $\text{Re}\,\tilde{h}_{\lambda}$ is logarithmic and dominates as $f\to 0$. These corrections are associated with the $1/t$ falloff of the waveform as $t\to\pm\infty$ (see Sec. II). As a measure of the relative deviation from the soft waveform, we show the contours of $|f\tilde{h}_{2,\pm 2}/(\lim_{f\to 0}f\tilde{h}_{2,\pm 2})-1|$ and $|f\tilde{h}_{2,0}/(\lim_{f\to 0}f\tilde{h}_{2,\pm 2})|$ in Fig. 1. Another measure is the deviation of $\text{arg}\,f\tilde{h}$ from its soft limit, for which a concrete example is shown in Fig. 2. From Eq. (21), the memory amplitude $|\Delta h|=\sqrt{\Delta h_{+}^{2}+\Delta h_{\times}^{2}}=\frac{\nu M}{R}\frac{\sqrt{e^{2}-1}}{ae^{2}}\sqrt{-8s^{4}_{\Theta}c_{4\Phi}+28c_{2\Theta}+c_{4\Theta}+35}$. Thus $|\Delta h|\propto M\sqrt{e^{2}-1}\,/(ae^{2})=(b/M)/[(b/M)^{2}+1/v_{0}^{4}]$, which increases with $v_{0}$ for given $b$, and peaks at $b=M/v_{0}^{2}$ for given $v_{0}$. For given $b$ and $v_{0}$, $|\Delta h|$ takes the maximum value $8$ at $\Theta\in\{0,\pi\}$, and vanishes at $\Theta=\pi/2\,\land\,\Phi\,\text{mod}\,\pi/2=0$.
@@
$H_{in}$ $\overset{n\to\infty}{\to}-\frac{ie^{-n\left(\sqrt{e^{2}-1}-\text{arcsec}\,e\right)}}{(e^{2}-1)^{1/4}\sqrt{n\pi/2}},$ (22)
$H_{1+in}$ $\overset{n\to\infty}{\to}-\frac{\left(\sqrt{e^{2}-1}+i\right)e^{-n\left(\sqrt{e^{2}-1}-\text{arcsec}\,e\right)}}{e(e^{2}-1)^{1/4}\sqrt{n\pi/2}}.$
@@
Figure 2: Relative deviations from the soft waveform in $\text{arg}\,f\tilde{h}$ for various eccentricities, at $\{\Theta,\Phi\}=\{\pi/6,\pi/3\}$. Dashed lines correspond to the negative-frequency spectrum, $f\,\tilde{h}(-f)$. Black dot-dashed lines show the logarithmic fit at $\varpi=\pi f/\Omega=10^{-5}$, $\text{arg}\,f\tilde{h}-\text{arg}\,f\tilde{h}(f\to 0)\propto\varpi\ln\varpi$.
Several features outlined above persist once PN corrections are included, at least when those corrections are not too large. We include the PN corrections to the CM-frame equation of motion (EOM) $\ddot{\mathbf{x}}=-(M/r^{3})\,\mathbf{x}+\mathbf{F}$ and the gravitational waveform according to [65]. The initial condition is set by an osculating hyperbolic orbit with orbital elements $\{a,e\}$ and the total mass parameter $M$ (see, e.g., [18]). Such a choice for the osculating orbit is not unique, since the initial orbit depends only on $\mathbf{v}_{0}$ and $\mathbf{b}$. However, this setting enable us to directly compare the results of Newtonian and PN EOMs in the harmonic coordinates. We choose the initial conditions such that the osculating Keplerian orbit coincides with the one considered above. For simplicity, we neglect the spins of the two bodies, thus $\mathbf{F}=\mathbf{F}_{\text{1PN}}+\mathbf{F}_{\text{2PN}}+\mathbf{F}_{\text{2.5PN}}+\mathcal{O}(\text{3PN})$; the PN orbit also lies in the $XY$ plane, but its periastron deviates from that of the Keplerian orbit. At 1PN order, the change of the argument of periastron $\varphi_{0}$ of the osculating orbit during the scattering is given by [18, 13]
@@
For concreteness, we choose $\nu=0.2$, $b/M=100$ and $v_{0}=0.13$, the results are shown in Fig. 3. To verify the convergence of PN corrections, we perform the computation at different PN orders. The approximation (23) (shown as the thick gray line in the top panel of Fig. 3) is found to agree well with the 2.5PN orbit, indicating that higher-order PN corrections are relatively small. The orbit is therefore well described by the 1PN-accurate quasi-Keplerian parameterization [26]. Nevertheless, there is a considerable difference between the 0PN and 0.5PN waveforms, but the 1PN correction to the waveform is already small. Note that the null memory contribution to the waveform in the present case enters at 2.5PN order [35], and is thus negligible.
@@
As a test, we verify that the spectrum of 0PN waveform computed using 0PN orbit agrees well with Eqs. (19)-(20). We also find that when the exact correction factor $\mathcal{C}(f_{\text{min}})$ is small, the linear fit performed near $f_{\text{min}}$ provides a reliable extrapolation to frequencies $0<f<f_{\text{min}}$. This extrapolation improves as $f_{\text{min}}$ decreases. The spectra of $|\text{Re}\,\tilde{h}_{\pm}(f)|$ and $|\text{Im}\,\tilde{h}_{\pm}(f)|$ of the 1PN waveforms are shown in the bottom panel of Fig. 3. The numerical spectrum at the lowest frequency is pretty close to the soft waveform with $\Delta h_{\lambda}$ computed in the time domain. The spectrum of $f\,\text{Im}\,\tilde{h}_{\pm}$ at low frequencies is well described by a linear correction: $f\,\text{Im}\,\tilde{h}_{\lambda}\propto 1+\kappa_{\lambda}\varpi$, with $\kappa_{+}\approx 0.95\pi$ and $\kappa_{\times}\approx 0.92\pi$. These slightly deviate from the Keplerian value $\kappa=\pi$ when expanded near $f=0$, as given by Eq. (21). In this example, at high frequencies, $\text{Im}\,\tilde{h}_{+}\approx\text{Re}\,\tilde{h}_{\times}$ and $\text{Re}\,\tilde{h}_{+}\approx\text{Im}\,\tilde{h}_{\times}$, indicating that the radiation is nearly right-handed polarized. [^6] Using (22), the high-frequency limit of the 0PN waveform is dominated by $|m|=2$ modes and becomes elliptically polarized, since $\lim_{f\to\infty}(\tilde{h}_{+},\tilde{h}_{\times})\propto e^{2i\Phi}(3+c_{2\Theta},4ic_{\Theta})$, similar to that of a circular orbit. This is also reflected in the 0PN spectra of the $Z$-component angular momentum and graviton number fluxes, whose ratio approaches 2 at $f\to\infty$ (as for an elliptical orbit). The decaying behavior of the spectrum at high frequencies is similar to the 0PN approximation given by Eq. (22), as depicted by the green lines in the bottom panel of Fig. 3.
@@
Fig. 4 shows the spectrum $\tilde{h}^{\prime}(f)$ obtained by a phase rotation that transforms ${i(\Delta h_{+}-i\Delta h_{\times})}/{(2\pi f)}$ to a purely imaginary value. The negative-frequency spectrum is obtained from the Fourier transform of $h^{*}(t)$ under the same phase rotation. We see that the positive-frequency spectrum of $\tilde{h}^{\prime}(f)$ can be described by a soft waveform with a correction factor $C(f)=1+\kappa\varpi+\mathcal{O}(\varpi^{2})$ at low frequencies. In this example, $\kappa\approx 2.63\pi-(0.19\pi)i$, which is quite different from $\kappa_{\lambda}$ describing the linear correction to $f\,\text{Im}\,\tilde{h}_{\lambda}$.
@@
Figure 3: CM-frame orbits (top panel) and gravitational waveforms in the time domain (middle panel) and frequency domain (bottom panel) of a compact binary scattering, with $\nu=0.2$, $b/M=100$, $v_{0}=0.13$ (corresponding to $e=2$), and $\{\Theta,\Phi\}=\{\pi/6,\pi/3\}$. In the middle panel, the 0PN waveform is computed using the 0PN orbit, whereas the 0.5PN and 1PN waveforms are computed using the 2PN orbit. The bottom panel shows the spectrum of the 1PN waveform, with $\Omega=\sqrt{M/a^{3}}=v_{0}^{3}/M\approx\pi(v_{0}/0.13)^{3}(100\,M_{\odot}/M)\times 1.4\,\text{Hz}$. Before computing the spectrum, we shift the time coordinate such that the periastron crossing happens at $t=0$. The horizontal line in the bottom panel corresponds to $\Delta h_{\lambda}$ of the time-domain waveform.
@@
Figure 4: Spectra of $f\,\text{Im}\,\tilde{h}^{\prime}(f)$ and $f\,\text{Re}\,\tilde{h}^{\prime}(f)$. Dashed lines denote the negative-frequency spectrum, $f\,\tilde{h}^{\prime}(-f)$. The linear and quadratic fits to the low-frequency part of the positive-frequency spectra are also shown. The horizontal line corresponds to $\Delta h$ of the time-domain waveform.
@@
For an extreme mass-ratio scattering involving a large BH, the orbit of the small compact body can be approximated by an unbound geodesic in the background BH spacetime. The dissipative effect is expected to be negligible in this case, since the incident velocity for scattering is bounded from below by the critical velocity for plunge in the test-body approximation. We compute the ordinary displacement memory produced by such a system using the general formula [12]:
@@
This is derived from the linearized Einstein equation in a flat background, but can be applied to scattering processes if the initial and final states live in a flat background. Eq. (24) can also be derived from the soft graviton theorem [85]. The low-velocity expansion of Eq. (24) up to $\mathcal{O}(v^{4})$ (using the 1PN approximation for $\mathbf{v}_{1,2}$ in the CM frame) agrees with the 1PN approximation:
@@
$\quad+(1-3\nu)\left[(\mathbf{v}\cdot\hat{\mathbf{k}})^{2}+v^{2}/2\right]\Big)\Big]\frac{2\nu M}{R}.$
For simplicity, we consider a non-spinning central BH, and thus compute the deflection angle $\chi$ of the small test body with mass $\nu M$ in a Schwarzschild spacetime with mass parameter $M$. The source frame is again chosen such that the initial orbit lies in the $XY$ plane, the initial and final three-velocities are along the direction $\mathbf{N}_{\text{in}}=c_{\vartheta}\mathbf{e}_{X}+s_{\vartheta}\mathbf{e}_{Y}$ and $\mathbf{N}_{\text{out}}=-c_{\vartheta}\mathbf{e}_{X}+s_{\vartheta}\mathbf{e}_{Y}$, respectively, with $\vartheta=(\chi-\pi)/2$. The displacement memory from Eq. (24) is given by
@@
Figure 5: Ordinary displacement memory from the scattering of a test body with mass $\nu M$ by a non-spinning BH with mass $M$, for $b/M=100$, and $\{\Theta,\Phi\}=\{\pi/5,\pi/7\}$. The blue and black solid lines show the test-body approximation. The red (green) dashed lines show the 0PN (1PN) approximations. The pink dashed lines show the results computed from Eq. (26), using the 1PN approximation to the deflection angle. The critical incident velocity $v_{\text{plunge}}\approx 0.04$. The figure inset shows the results for small incident velocities, with the horizontal axis given by $v_{0}-v_{\text{plunge}}$. The gray dot-dashed lines show the results for a central BH with dimensionless spin $\chi=1$ along the $Z$-axis, for which $v_{\text{plunge}}=0.02$.
@@
The deflection angle depends on the initial velocity $v_{0}$ and the dimensionless impact parameter $b/M$. For the small body to be scattered rather than plunge into the BH, the impact parameter needs to exceed a critical value: ${b_{\text{plunge}}}/{M}=\sqrt{4+\frac{\sqrt{8v_{0}^{2}+1}-1}{2v_{0}^{4}}+\frac{4\sqrt{8v_{0}^{2}+1}+10}{v_{0}^{2}}}$. For a null particle, $b_{\text{plunge}}(v_{0}=1)=3\sqrt{3}M$. Likewise for a given impact parameter, $v_{0}$ needs to exceed a critical value $v_{\text{plunge}}$ in order to escape from the BH. It is possible for high-energy massless or ultra-relativistic particles to be scattered by a BH, but this cannot produce an observable memory, since even a photon with $\omega=10\,\text{TeV}$ amounts to a mass of only $9\times 10^{-54}\,M_{\odot}$.
@@
For the scattering of timelike particles, a concrete example is shown in Fig. 5, where the results are compared with the 0PN approximation in Eq. (21) (red dashed lines) and the results computed using Eq. (25) and the 1PN approximation to the deflection angle: $\chi=2\arccos(-1/e)-\pi+\Delta\varphi_{0}$ (green dashed lines), with $\Delta\varphi_{0}$ given by Eq. (23). For comparison, we also show the results (pink dashed lines) computed from Eq. (26), but using the 1PN approximation to the deflection angle, which remains in good agreement with the test-body approximation even at high velocities. As can be seen, the 0PN approximation is accurate only for intermediate values of $v_{0}$, and the situation near $v_{\text{plunge}}$ cannot be captured by the 1PN deflection angle either, where the produced memory is relatively small. For $v_{0}\gtrsim v_{\text{plunge}}$, the periastron occurs at $r\gtrsim 4M$; the result is therefore sensitive to the spin of the central BH. Eq. (26) can also be applied to the scattering orbit in the ecliptic plane of Kerr spacetime; the gray dot-dashed lines in Fig. 5 show the result for a central BH with dimensionless spin $\chi=1$ along the $Z$-axis, which are, interestingly, closer to the pink dashed lines. At high velocities, the amplitudes of memory are considerably larger than the 0PN prediction. As in the 0PN approximation, $|\Delta h|$ is invariant under the transformation $\Theta\to\pi-\Theta$ or $\Phi\to\pi-\Phi$, but its extreme point deviates from $\Theta\in\{0,\pi\}$. For $\Theta=\pi/2$, $|\Delta h|$ vanishes at four values of $\Phi$, two of which occur at $\pi/2$ and $3\pi/2$, while the remaining two are displaced from 0 and $\pi$.
III.3 Displacement memory from BBH merger